A 8.75- g bullet from a 9-mm pistol has a velocity of 397.0 m/s. It strikes the 0.705- kg block of a ballistic pendulum and passes completely through the block. If the block rises through a distance h = 21.83 cm, what was the velocity of the bullet as it emerged from the block?


1
Expert's answer
2021-03-29T09:00:02-0400

"\text{System energy before bullet impact:}"

"E_1 =\frac{mV_1^2}{2}+Mgh_1;"

"\text{System energy after bullet impact:}"

"E_2 =\frac{mV_2^2}{2}+Mgh_2;"

"\text{where:}"

"m=8.75*10^{-3}kg-\text{bullet mass}"

"M=0.705 \ kg\text{ballistic pendulum mass}"

"V_1,V_2\text{ bullet velocity before and after impact}"

"h_1,h_2 \text{ ballistic pendulum height above ground }h_2-h_1=0.2183\ m"

"E_1=E_2"

"\frac{mV_1^2}{2}+Mgh_1=\frac{mV_2^2}{2}+Mgh_2"

"\frac{mV_1^2}{2}-\frac{mV_2^2}{2}=Mgh_2-Mgh_1"

"\frac{mV_2^2}{2}=\frac{mV_1^2}{2}-Mg(h_2-h_1)"

"\frac{mV_2^2}{2}="

"V_2=\sqrt{(V_1^2-\frac{2Mg(h_2-h_1)}{m})}\approx396.56"

Answer:396.56m/s